# On the hypothesis of a bi-isotropic plasma permeating the interstellar space

### v1

@@ Abstract
In this work, we study the propagation of electromagnetic waves in a magnetized chiral plasma that pervades the interstellar space. The Maxwell equations, supplemented by bi-isotropic constitutive relations, are rewritten to describe a cold, uniform, and collisionless plasma model that yields new collective electromagnetic modes for distinct pairs of refractive indices associated with right- and left-handed circularly polarized waves. We have investigated the optical behavior through the rotatory power (RP) and dichroism coefficient, reporting that the finite chiral parameter induces double RP sign reversal, an exotic optical signature that takes place in chiral dielectrics and rotating plasmas. In the low-frequency regime, a modified propagating helicon with right-handed circular polarization is obtained. Next, supposing that the interstellar medium behaves as a bi-isotropic cold plasma, we employ Astrophysical data of radio pulsars to achieve upper limits on the magnetoelectric parameters magnitude. In particular, by using dispersion measure and rotation measure data from five pulsars, we constrain the magnitude of the chiral parameter to the order of $10^{-16}$ and $10^{-22}$, respectively.
@@ I Introduction
Chiral optically active media are characterized by the absence of inversion symmetry (parity), being described by parity-odd electromagnetic models [1, 2, 3]. Optical activity makes left- and right-handed circularly polarized (LCP and RCP) waves propagate with distinct phase velocities, implying birefringence [4], phenomenon that may arise from natural features of the medium or can be induced by external fields (e.g., Faraday effect [5, 6, 7]). Birefringence is generally quantified in terms of the rotation angle per unit length or rotatory power (RP) [8], providing a useful way to probe non-usual electromagnetic responses [9, 10] and also to discuss underlying physics in context of particle physics and cosmology [11]. To account for such an effect in chiral systems and other optical-related phenomena, extensions of the standard isotropic description of the medium’s electromagnetic response are often used [12, 13, 14, 15]. For instance, in recent decades, bi-isotropic [16, 17] and bi-anisotropic electrodynamics [22, 18, 19, 20, 21] have been considered to describe optically active systems, whose constitutive relations can be generally written as
@@
Bi-isotropic constitutive relations are widely employed in modified electromagnetic field theories and condensed matter physics to describe optical properties of materials [24, 25, 26], the electromagnetic response of topological insulators [27, 28, 29, 31, 32, 33, 30], axion couplings [34, 35, 36, 37], Casimir effect in chiral media [38, 39], and applications in photonic systems [40]. Bi-anisotropic magnetoelectric parameters are generally encoded in the off-diagonal elements of the matrices $\alpha_{ij}$, $\beta_{ij}$ of relations (1a) and (1b), being used for investigating surface plasmon polaritons on bianisotropic substrates [41], group and energy velocities waves in heterostructures and surfaces [42], wave propagation in time-dependent media with antisymmetric magnetoelectric coupling [43], Weyl semimetals [44, 45], ferromagnetic gyromagnetic materials [46], the discovery of electromagnons in perovskites associated with giant directional dichroism [47], and enhanced gyrotropic birefringence in a helimagnet [48]. Anisotropic magnetoelectric coefficients are important to describe magneto-optical effects of topological insulators [49, 50] and graphene compounds [51].
@@
is obtained for a magnetic background $\mathbf{B}=B_{0}\hat{z}$. Here, $\epsilon_{ijk}$ is the Levi-Civita symbol with $\epsilon_{123}=1$, $S=1-{\omega_{p}^{2}}/{(\omega^{2}-\omega_{c}^{2})}$, $D={\omega_{c}\omega_{p}^{2}}/({\omega(\omega^{2}-\omega_{c}^{2})})$, and $P=1-{\omega_{p}^{2}}/{\omega^{2}},$ in which $\omega_{p}^{2}=nq^{2}/(\varepsilon_{0}m)$ and $\omega_{c}=qB_{0}/m$ are the squared plasma and cyclotron frequencies for particles of mass $m$ and charge $q$, respectively. This system yields the well-known longitudinal plasma oscillation at $\omega=\omega_{p}$, and the transversal left- and right-handed circularly polarized modes, associated respectively with the refractive indices
@@
Plasma physics is closely related to two relevant astrophysical observables used for examining dispersion and polarization rotation of waves coming from distant sources: the dispersion measure (DM) and rotation measure (RM). For radio waves travelling a distance $d$ through the interstellar medium (ISM) described by a cold, ionized bi-isotropic chiral plasma, one defines the arrival time, $t=\int_{0}^{d}(v_{g})^{-1}ds$, where $s$ is the line of sight element and $v_{g}$ the group velocity [55]. For frequencies large compared to the plasma frequency, i.e, $\omega\gg\omega_{p}$, the ISM does not present absorption (lossless media). In this range, the usual plasma refractive indices [^1] We consider an electron plasma, where the cyclotron frequency $\omega_{c}$ incorporates the electron charge q=eq=-e. (4) provide
$v_{g}^{-1}\approx 1+\frac{\omega_{p}^{2}}{2\omega^{2}}\pm\frac{\omega_{c}\omega_{p}^{2}}{\omega^{3}}.$ (5)
In the latter, the term in $\omega^{-3}$ is smaller than the second one by the factor $\omega_{c}/\omega\sim 10^{-7}$ (for an interstellar magnetic field of the order of a few $\mu\mathrm{G}$), so that it can be neglected, yielding the following arrival time:
$t\approx d+\frac{e^{2}}{2m\omega^{2}}\text{DM},\quad\text{DM}=\int^{d}_{0}n_{e}ds.$ (6)
@@
$\tau=\frac{e^{2}}{2m\omega^{2}}\text{DM},$ (7)
@@
$\Delta\Psi=\int_{0}^{d}\left(k_{R}-k_{L}\right)ds=\frac{e^{3}\lambda^{2}}{4\pi^{2}m^{2}}\int_{0}^{d}n_{e}B_{\parallel}ds,$ (8)
@@
$k_{R,L}\approx\omega-\frac{\omega_{p}^{2}}{2\omega}\pm\frac{\omega_{c}\omega_{p}^{2}}{2\omega^{2}}.$ (9)
The third term is smaller than the second one by the factor $\omega_{c}/\omega\sim 10^{-7}$. However, it is the leading order term for the phase shift of Eq. (8) and determines the Faraday effect under the small density condition. Thus, the differential phase rotation (8) yields the polarization rotation angle
@@
$\text{RM}=\frac{e^{3}}{8\pi^{2}m^{2}}\int_{0}^{d}n_{e}B_{\parallel}ds,$ (11)
@@
given in $\mathrm{rad}/\mathrm{m}^{2}$. From the electromagnetic time delay, relevant information to estimate the Galactic electron distribution permeating the ISM [57, 56, 55] are collected, in which the DM is a key parameter for studying dispersive ISM effects along the wave path [58]. On the other hand, the rotation measure (RM) is a widely employed tool for estimating the magnitude and direction of the Galactic magnetic fields, derived from analyzing the polarization of light emitted by pulsars [59, 60, 61].
Observations of radio pulsars by multiple telescopes have provided datasets that include several relevant quantities of these objects, such as DM and RM. The LOFAR collaboration [62] has published data on DMs, flux densities, and calibrated total intensity profiles for a subset of pulsars observed by its high-band antennas (110–188 MHz). Additionally, LOFAR has significantly enhanced the precision of RM measurements for low-frequency pulsars [63].
@@
As is well known, bi-isotropic electrodynamics constitutes an effective tool to describe real phenomena optical properties and wave propagation in chiral matter. On the other hand, in this work we examine the possibility of the plasma being described by such extended constitutive relations. We develop such an investigation analyzing optical properties of a magnetized cold plasma ruled by bi-isotropic constitutive relations (12). In Sec. II, we determine the modified refractive indices and analyze the propagation and absorption zones. In Sec. III, we discuss optical effects of the chiral plasma, such as birefringence and dichroism, supposing a finite magnitude chiral factor. In Sec. IV, astrophysical data of DM and RM of five pulsars are employed to constrain the magnitude of the chiral parameter, assuming that radio waves travel in an interstellar bi-isotropic medium (ISM). Finally, we summarize the results and perspectives in Sec. V.
II Wave propagation in bi-isotropic chiral plasma
@@
Chiral cold plasmas are another type of system exhibiting optical activity and have been investigated in at least two distinct scenarios: i) in the axion-like Maxwell-Carroll-Field-Jackiw (MCFJ) electrodynamics [64, 65], considering the influence of the chiral magnetic current and anomalous Hall current on the propagating modes and optical features [66, 67, 68]; ii) in bi-isotropic Maxwell electrodynamics [70, 69, 71], where negative refraction emerges as a consequence of the chiral parameter. In this latter case, the bi-isotropic constitutive relations are
$\mathbf{D}$ $=\hat{\epsilon}\mathbf{E}+i\xi_{c}\mathbf{B},$ (12a)
$\mathbf{H}$ $=i\xi_{c}\mathbf{E}+\mu^{-1}\mathbf{B},$ (12b)
which hold under the condition (2).
Using Eq. (12) in Maxwell equations yields the modified wave equation
@@
whose non-trivial solutions arise from the condition that the determinant of the coefficient matrix vanishes, yielding
@@
  • The refractive index $n_{R_{+}}$, illustrated in Fig. 1, presents a free propagation (without absorption) zone for $0<\omega<\omega_{c}$, similarly to the usual case (solid black line). For $\omega_{c}<\omega<\omega_{R-}$, the index is complex, $\mathrm{Im}[n_{R_{+}}]\neq 0$, and an absorption zone occurs (where the real part of the index is negative). As $\omega_{R-}<\omega_{-}$, one has a shortened absorption zone compared to the usual case, see Fig. 1. In addition, a free propagating negative refraction zone (non-existent in the standard case) emerges for $\omega_{R-}<\omega<\omega_{-}$.

### v2

# Hypothesis of a bi-isotropic-like plasma permeating the interstellar space

@@ Abstract
In this work, we study the propagation of electromagnetic waves in a magnetized chiral plasma that pervades the interstellar space. The Maxwell equations, supplemented by bi-isotropic-like constitutive relations, are rewritten to describe a cold, uniform, and collisionless plasma model that yields new collective electromagnetic modes for distinct pairs of refractive indices associated with right- and left-handed circularly polarized waves. We have investigated the optical behavior through the rotatory power (RP) and dichroism coefficient, reporting that the finite chiral parameter induces double RP sign reversal, an exotic optical signature that takes place in chiral dielectrics and rotating plasmas. In the low-frequency regime, a modified propagating helicon with right-handed circular polarization is obtained. Next, supposing that the interstellar medium behaves as a chiral bi-isotropic-like cold plasma, we employ Astrophysical data of radio pulsars to achieve upper limits on the magnetoelectric parameters magnitude. In particular, by using dispersion measure and rotation measure data from five pulsars, we constrain the magnitude of the chiral parameter to the order of $10^{-16}$ and $10^{-22}$, respectively.
@@ I Introduction
Chiral optically active media are characterized by the absence of inversion symmetry (parity), being described by parity-odd electromagnetic models [1, 2, 3]. Optical activity makes left- and right-handed circularly polarized (LCP and RCP) waves propagate with distinct phase velocities, implying birefringence [4], phenomenon that may arise from natural features of the medium or can be induced by external fields (e.g., Faraday effect [5, 6, 7]). Birefringence is generally quantified in terms of the rotation angle per unit length or rotatory power (RP) [8], providing a useful way to probe nonusual electromagnetic responses [9, 10] and also to discuss underlying physics in context of particle physics and cosmology [11]. To account for such an effect in chiral systems and other optical-related phenomena, extensions of the standard isotropic description of the medium’s electromagnetic response are often used [12, 13, 14, 15]. For instance, in recent decades, bi-isotropic [16, 17] and bianisotropic electrodynamics [22, 18, 19, 20, 21] have been considered to describe optically active systems, whose constitutive relations can be generally written as
@@
Bi-isotropic constitutive relations are widely employed in modified electromagnetic field theories and condensed matter physics to describe optical properties of materials [24, 25, 26], the electromagnetic response of topological insulators [27, 28, 29, 31, 32, 33, 30], axion couplings [34, 35, 36, 37], Casimir effect in chiral media [38, 39], and applications in photonic systems [40]. Bi-anisotropic magnetoelectric parameters are generally encoded in the off-diagonal elements of the matrices $\alpha_{ij}$, $\beta_{ij}$ in relations (1a) and (1b), being used for investigating surface plasmon polaritons on bianisotropic substrates [41], group and energy velocities waves in heterostructures and surfaces [42], wave propagation in time-dependent media with antisymmetric magnetoelectric coupling [43], Weyl semimetals [44, 45], ferromagnetic gyromagnetic materials [46], discovery of electromagnons in perovskites associated with giant directional dichroism [47], and enhanced gyrotropic birefringence in a helimagnet [48]. Anisotropic magnetoelectric coefficients are also important to describe magneto-optical effects of topological insulators [49, 50] and graphene compounds [51].
@@
is obtained for a magnetic background, $\mathbf{B}=B_{0}\hat{z}$. Here, $\epsilon_{ijk}$ is the Levi-Civita symbol with $\epsilon_{123}=1$, $S=1-{\omega_{p}^{2}}/{(\omega^{2}-\omega_{c}^{2})}$, $D={\omega_{c}\omega_{p}^{2}}/({\omega(\omega^{2}-\omega_{c}^{2})})$, and $P=1-{\omega_{p}^{2}}/{\omega^{2}},$ in which $\omega_{p}^{2}=nq^{2}/(\varepsilon_{0}m)$ and $\omega_{c}=qB_{0}/m$ are the squared plasma and cyclotron frequencies for particles of mass $m$ and charge $q$, respectively. This system yields the well-known longitudinal plasma oscillation at $\omega=\omega_{p}$, and the transversal left- and right-handed circularly polarized modes, associated respectively with the refractive indices
@@
Plasma physics is closely related to two relevant astrophysical observables used for examining dispersion and polarization rotation of waves coming from distant sources: the dispersion measure (DM) and rotation measure (RM). For radio waves traveling a distance $d$ through the interstellar medium (ISM) described by a cold, ionized bi-isotropic chiral plasma, one defines the arrival time, $t=\int_{0}^{d}(v_{g})^{-1}ds$, where $s$ is the line of sight element and $v_{g}$ the group velocity [55]. For frequencies large compared to the plasma frequency, i.e, $\omega\gg\omega_{p}$, the ISM does not present absorption (lossless media). In this range, the usual plasma refractive indices [^1] We consider an electron plasma, where the cyclotron frequency $\omega_{c}$ incorporates the electron charge q=eq=-e. (4) provide
$v_{g}^{-1}\approx\frac{1}{c}+\frac{\omega_{p}^{2}}{2c\omega^{2}}\pm\frac{\omega_{c}\omega_{p}^{2}}{c\omega^{3}},$ (5)
whose term in $\omega^{-3}$ is smaller than the second one by the factor $\omega_{c}/\omega\sim 10^{-7}$ (for an interstellar magnetic field of the order of a few $\mu\mathrm{G}$). Thus, it can be neglected, yielding the following arrival time
$t\approx\frac{d}{c}+\frac{e^{2}}{2c\epsilon_{0}m\omega^{2}}\text{DM},\quad\text{DM}=\int^{d}_{0}n_{e}ds.$ (6)
@@
$\tau=\frac{e^{2}}{2c\epsilon_{0}m\omega^{2}}\text{DM},$ (7)
@@
$\Delta\Psi=\int_{0}^{d}\left(k_{R}-k_{L}\right)ds=\frac{e^{3}\lambda^{2}}{4\pi^{2}c^{3}m^{2}\epsilon_{0}}\int_{0}^{d}n_{e}B_{\parallel}ds,$ (8)
@@
$k_{R,L}\approx\frac{\omega}{c}-\frac{\omega_{p}^{2}}{2c\omega}\pm\frac{\omega_{c}\omega_{p}^{2}}{2c\omega^{2}}.$ (9)
The third term, despite smaller than the second one, is the leading order term for the phase shift of Eq. (8) and determines the Faraday effect under the small density condition. Thus, the differential phase rotation (8) yields the polarization rotation angle
@@
$\text{RM}=\frac{e^{3}}{8\pi^{2}c^{3}m^{2}\epsilon_{0}}\int_{0}^{d}n_{e}B_{\parallel}ds,$ (11)
@@
given in $\mathrm{rad}/\mathrm{m}^{2}$. From the electromagnetic time delay, relevant information to estimate the Galactic electron distribution permeating the ISM [57, 56, 55] are collected, in which the DM is a key parameter for studying dispersive ISM effects along the wave path [58]. On the other hand, the RM is a widely employed tool for estimating the magnitude and direction of the Galactic magnetic fields, derived from analyzing the polarization of light emitted by pulsars [59, 60, 61]. Observations of radio pulsars by multiple telescopes have provided datasets that include several relevant quantities of these objects, such as DM and RM. The LOFAR collaboration [62] has published data on DMs, flux densities, and calibrated total intensity profiles for a subset of pulsars observed by its high-band antennas (110–188 MHz). Additionally, LOFAR has significantly enhanced the precision of RM measurements for low-frequency pulsars [63]. The effects of a chiral plasma permeating the ISM, on the DM and RM of light coming from pulsars, were recently investigated in Ref. [64], yielding upper bounds on the chiral factors.
One way of engendering a chiral plasma is to consider the chiral magnetic effect (CME) [65, 66], a phenomenon arising from an imbalance between right- and left-handed fermions which induces an electrical current proportional to the magnetic field, $\mathbf{J}=\sigma_{B}\mathbf{B}$, where $\sigma_{B}$ represents the magnetic conductivity. The emergence of such an effect leads to chiral plasma instabilities [67, 68, 70, 69] and interesting astrophysical repercussions [71]. Indeed, the CME plays an important role in the amplification and evolution of magnetic fields in the early Universe [72, 73], in which the cosmic medium behaves as a chiral plasma. The chiral magnetic instability was further investigated in a protoneutron star during its early cooling phase (after the supernova core collapse), where it contributes to generate magnetic fields as large as $10^{14}$ G [74]. The CME can also occur in the magnetosphere of magnetars, triggering chiral plasma instability that can be associated with the emission of circularly polarized electromagnetic radiation in a wide window of frequencies, possibly affecting signals from pulsars and fast radio bursts features [75]. A recent investigation has shown that the chiral magnetic current acts analogously to a dynamo, amplifying magnetic fields by using the energy stored in the chiral chemical potential, providing toroidal and poloidal large-scale dipolar magnetic components of $10^{14}$ G [76]. Furthermore, parity-violating effects may emerge from a chiral primordial gravitational-wave background [77, 78, 79, 80, 81, 82] that would induce chiral photons through photon-graviton interaction [83], contributing to the CMB polarization [84].
The chiral magnetic current can be effectively described within the framework of the CPT-odd Maxwell-Carroll-Field-Jackiw electrodynamics (MCFJ) [85, 86, 87], a Lorentz-violating extension of Maxwell theory whose Lagrangian [^3] Here, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ and $G^{\mu\nu}=\frac{1}{2}\chi^{\mu\nu\alpha\beta}F_{\alpha\beta}$ are the usual $U(1)$ vacuum and continuous matter field strength, respectively, where $\chi^{\mu\nu\alpha\beta}$ describes the medium constitutive tensor.
$\mathrm{{\mathcal{L}}}=-\frac{1}{4}G^{\mu\nu}F_{\mu\nu}+\frac{1}{4}\epsilon^{\mu\nu\alpha\beta}\left(k_{AF}\right)_{\mu}A_{\nu}F_{\alpha\beta},$ (12)
yields the inhomogeneous equations
$\nabla\cdot\mathbf{D}+\mathbf{k}_{AF}\cdot\mathbf{B}$ $=0,$ (13)
$\nabla\times\mathbf{H}-\frac{\partial\mathbf{D}}{\partial t}$ $=-k_{AF}^{0}\mathbf{B}+\mathbf{k}_{AF}\times\mathbf{E}.$ (14)
The modified Ampere’s law incorporates anomalous current terms, where the factor $k_{AF}^{0}$ describes the magnetic conductivity and the vector $\mathbf{k}_{AF}$ represents the anomalous Hall conductivity. For the vacuum case, one writes
$\epsilon_{0}\nabla\cdot\mathbf{E}$ $=-\mathbf{k}_{AF}\cdot\mathbf{B},$ (15)
$\frac{1}{\mu_{0}}\nabla\times\mathbf{B}-\epsilon_{0}\frac{\partial\mathbf{E}}{\partial t}$ $=-k_{AF}^{0}\mathbf{B}+\mathbf{k}_{AF}\times\mathbf{E}.$ (16)
The CFJ Lagrangian can also be obtained from the axion coupling,
$\mathrm{{\mathcal{L}}}_{axion}=-\frac{1}{4}F^{\mu\nu}F_{\mu\nu}+\theta\left(\mathbf{r},t\right)\left(\mathbf{E}\cdot\mathbf{B}\right),$ (17)
by means of an integration, which reveals the connection $\partial_{\mu}\theta=-\left(k_{AF}\right)_{\mu}$ [34]. The axion electrodynamics can also be addressed by introducing bi-isotropic constitutive relations [88, 89, 90],
$\mathbf{D}$ $=\epsilon_{0}\mathbf{E}+\theta\mathbf{B},$ (18)
$\mathbf{H}$ $=\mu_{0}^{-1}\mathbf{B}-\theta\mathbf{E},$ (19)
into the Maxwell’s equations. In fact, by replacing the constitutive relation (18) in the Gauss’s law, $\nabla\cdot\mathbf{D}=\rho,$ one obtains
$\epsilon_{0}\nabla\cdot\mathbf{E}+(\nabla\theta)\cdot\mathbf{B}=\rho,$ (20)
since $\nabla\cdot\mathbf{B=0}$. In the absence of sources, $\rho=0,$ the latter equation recovers Eq. (15), with $\nabla\theta=\mathbf{k}_{AF}.$ At the same time, replacing the constitutive relations (18) and (19) in the usual Ampere’s law, yields
$\frac{1}{\mu_{0}}\nabla\times\mathbf{B}-\epsilon_{0}\frac{\partial\mathbf{E}}{\partial t}=\nabla\theta\times\mathbf{E}+\left(\frac{\partial\theta}{\partial t}\right)\mathbf{B}+\mathbf{j},$ (21)
which is equivalent to Eq. (16) with $\partial_{t}\theta=-\left(k_{AF}\right)_{0}$, $\nabla\theta=\mathbf{k}_{AF}$, and $\mathbf{j}=0$. These conditions suggest a possible equivalence between the MCFJ model and the bi-isotropic electrodynamics of Eqs. (18) and (19). A mapping between the magnetoelectric parameter and the magnetic conductivity (or CFJ timelike factor) is given
$\theta=i\left(k_{AF}\right)_{0}/\omega.$ (22)
The CFJ theory with $\left(k_{AF}\right)_{0}\neq 0$ provides an electromagnetic framework that describes plasmas with CME [91, 92], phenomenon with astrophysical applications/effects, which can be connected with the bi-isotropic plasma addressed in the present work, given the equivalence stated in Eq. (22).
@@
As is well known, bi-isotropic electrodynamics constitutes an effective tool to describe real phenomena optical properties and wave propagation in chiral matter. On the other hand, in this work we examine the possibility of the plasma being described by such extended constitutive relations. We develop such an investigation analyzing optical properties of a magnetized cold plasma ruled by bi-isotropic-like constitutive relations (23). In Sec. II, we determine the modified refractive indices and analyze the propagation and absorption zones. In Sec. III, we discuss optical effects of the chiral plasma, such as birefringence and dichroism, supposing a finite magnitude chiral factor. In Sec. IV, astrophysical data of DM and RM of five pulsars are employed to constrain the magnitude of the chiral parameter, assuming that radio waves travel in an interstellar bi-isotropic-like medium (ISM). Finally, we summarize the results and perspectives in Sec. V.
II Wave propagation in bi-isotropic-like chiral plasma
@@
Chiral cold plasmas are another type of system exhibiting optical activity and have been investigated in at least two distinct scenarios: i) in the axionlike MCFJ electrodynamics [85, 86], considering the influence of the chiral magnetic current and anomalous Hall current on the propagating modes and optical features [91, 92, 93]; ii) in bi-isotropic-like Maxwell electrodynamics [95, 94, 96], where negative refraction emerges as a consequence of the chiral parameter. In this latter case, the electric permittivity is given the tensor of Eq. (3) and the constitutive relations are
$\mathbf{D}^{i}$ $=\epsilon_{ij}(\omega)\mathbf{E}^{j}+i\xi_{c}\mathbf{B}^{i},$ (23a)
$\mathbf{H}^{i}$ $=i\xi_{c}\mathbf{E}^{i}+\mu_{0}^{-1}\mathbf{B}^{i},$ (23b)
which hold under the condition (2). Notice that replacing Eq. (23) in the usual Maxwell equations, one finds
$\frac{1}{\mu_{0}}(\nabla\times{\bf{B}})^{i}-\epsilon_{ij}(\omega)\partial_{t}{\bf{E}}^{j}$ $=2i\xi_{c}\partial_{t}{\bf{B}}^{i},$ (24)
which allows us to establish a connection between $\xi_{c}$ and the CFJ timelike factor as
$2\xi_{c}\omega=-k_{AF}^{0}.$ (25)
A similar correspondence has been pointed in Ref. [97]. Considering Eq. (22) and Eq. (25), one can interpret that the CFJ timelike factor can effectively describe a dispersive chiral electromagnetic response of bi-isotropic systems.
In the present work, we examine a chiral plasma described by the bi-isotropic-like constitutive relations of Eq. (23a) and Eq. (23b), that replaced in the Maxwell equations, yields the modified wave equation
@@
whose nontrivial solutions arise from the condition that the determinant of the coefficient matrix vanishes, yielding
@@
  • The refractive index $n_{R_{+}}$, illustrated in Fig. 1, presents a free propagation (without absorption) zone for $0<\omega<\omega_{c}$, similarly to the usual case (solid black line). For $\omega_{c}<\omega<\omega_{R-}$, the index is complex, $\mathrm{Im}[n_{R_{+}}]\neq 0$, and an absorption zone occurs (where the real part of the index is negative). As $\omega_{R-}<\omega_{-}$, one has a shortened absorption zone compared to the usual case, see Fig. 1. In addition, a free propagating negative refraction zone (nonexistent in the standard case) emerges for $\omega_{R-}<\omega<\omega_{-}$.